# Towards Robust Real-World Multivariate Time Series Forecasting: A Unified Framework for Dependency, Asynchrony, and Missingness

### v1

@@ Appendix F Forecast Visualization
Figure 10(a) illustrates the forecasts of four models including our CTF for channel 3 in the ETT1 dataset under Scenario 1 in Section 5. The prediction horizon shows periodic sharp drops and recoveries, which are critical for accurate forecasting. Our model closely tracks these rapid transitions, while baseline models tend to underpredict their magnitude. We attribute this to the interpolation-free nature of our model, which avoids the frequency biases introduced by linear interpolation.
We also examine Scenario 2 in Section 5, which involves both asynchronous sampling and block-wise test-time missing. Figure 10(b) presents a visualization of model predictions for the solar power channel from the SolarWind dataset. In this sample, a missing block of length 144 is present in the middle of the test input. This causes TimeMixer and TimeXer to produce noticeably noisy forecasts. Although CrossGNN shows relatively higher robustness, its predictions exhibit a drift in the overall trend and an underestimation of the periodic transitions. In contrast, our model accurately identifies the missing region and applies masking, effectively preventing the injection of unreliable inputs. As a result, it successfully captures both the timing and magnitude of the periodic transitions, even across the missing block.
@@ Appendix G Discussion
To jointly capture temporal dynamics and cross-channel dependencies under the practical settings, we proposed a unified attention masking strategy that merges intra- and inter-channel attention into a single attention layer. This design avoids the need for separate attention modules, reducing the number of parameters and simplifying the overall model architecture. This parameter efficiency is particularly beneficial for datasets with limited data, as it reduces the risk of overfitting while maintaining the model’s ability to capture essential temporal and cross-channel patterns. However, this unified structure introduces scalability challenges when applied to high-dimensional datasets. As the number of channels increases, the attention layer must process a quadratic number of token interactions within a single computation step, leading to substantial computational overhead and slower training. A promising direction to mitigate this issue is to optimize the attention computation, for example, by decomposing the attention map or distributing the computation across multiple smaller attention modules.

### v2

@@ Appendix F Forecast Visualization
Figure 10(a) illustrates the forecasts of four models including our CTF on channel 3 of the ETT1 dataset under Case 1 in Section 5. The prediction horizon shows periodic sharp drops and recoveries, which are critical for accurate forecasting. Our model closely tracks these rapid transitions, while baseline models tend to underpredict their magnitude. We attribute this to the interpolation-free nature of our model, which avoids the frequency biases introduced by linear interpolation.
We also examine Case 2 in Section 5, which involves both asynchronous sampling and block-wise test time missing. Figure 10(b) presents a visualization of model predictions for the solar power channel from the SolarWind dataset. In this sample, a missing block of length 144 is present in the middle of the test input. This causes TimeMixer++ and TimeXer to produce noticeably noisy forecasts. Although CrossGNN shows relatively higher robustness, its predictions exhibit an underestimation of the periodic transitions. In contrast, our model accurately identifies the missing region and applies masking, effectively preventing the injection of unreliable inputs. As a result, it successfully captures both the timing and magnitude of the periodic transitions, even across the missing block.
@@ Appendix G Discussion
To jointly capture temporal dynamics and cross-channel dependencies under practical settings, we proposed a unified attention masking strategy that merges intra- and inter-channel attention into a single attention layer. This design avoids the need for separate attention modules, reducing the number of parameters and simplifying the overall model architecture. This parameter efficiency is particularly beneficial for datasets with limited data, as it reduces the risk of overfitting while maintaining the model’s ability to capture essential temporal and cross-channel patterns. However, this unified structure introduces scalability challenges when applied to high-dimensional datasets. As the number of channels increases, the attention layer must process a quadratic number of token interactions within a single computation step, leading to substantial computational overhead and slower training. A promising direction to mitigate this issue is to optimize the attention computation, for example, by decomposing the attention map or distributing the computation across multiple smaller attention modules.
@@
Use of Large Language Models (LLMs) in Paper Writing
LLMs were used exclusively for light copy-editing (grammar, clarity, concision) and small LaTeX phrasing/formatting suggestions. We did not employ LLMs for retrieval or discovery (e.g., related-work search) or for research ideation. LLMs did not draft technical content or citations and did not contribute at a level comparable to a coauthor. All text and claims were written, checked, and finalized by the authors; any LLM-suggested edits were adopted only after manual verification to prevent hallucinations or unsupported statements.
Ethics Statement
This work focuses on methodological contributions for multivariate time-series forecasting. Our experiments rely primarily on publicly available datasets from established academic and institutional sources. In addition, we include one private dataset provided under confidentiality agreements. This dataset contains no personally identifiable information, and we followed all necessary protocols to ensure secure handling, access control, and compliance with data-sharing restrictions. No raw private data will be released; only aggregated statistics and trained model weights are reported in the paper.
Our proposed model, CTF, is designed as a general forecasting framework and does not target specific individuals, groups, or sensitive applications. Potential downstream uses of time-series forecasting models include socially beneficial applications (e.g., energy demand planning, environmental monitoring) as well as sensitive ones (e.g., financial or policy decision-making). We encourage responsible usage, with attention to fairness, transparency, and possible societal impacts. All experiments were conducted under standard computational settings, with environmental costs comparable to typical machine learning research practice.
Reproducibility Statement
We have taken multiple steps to ensure the reproducibility of our work. All model architectures, training procedures, and evaluation protocols are described in detail in the main text and Appendix. Hyperparameters, preprocessing steps, optimization settings, and model configurations are provided in Appendix A.2. For publicly available datasets, we specify their sources in Appendix A.1 to facilitate exact replication. For the private dataset used under confidentiality agreements, raw data cannot be released, but we provide aggregate statistics, evaluation splits, and implementation details sufficient for reproducing the reported results.
Our codebase (including model implementation, training scripts, and evaluation pipeline) will be made available upon publication. All experiments were conducted on standard hardware (NVIDIA RTX 3090 GPUs) with fixed random seeds to reduce variance; each experiment was repeated across three random seeds, and average results are reported. We also provide details on software dependencies (Python version, PyTorch version, and major libraries) to ensure compatibility with common research environments.